OpenStudy (anonymous):

determine graphically the x-value when y=-5 in -x+2y=-6

OpenStudy (jackellyn):

Plug in y value into the equation and it will give you x, plot (x,y) on a graph

OpenStudy (anonymous):

-x+-10=-6?

OpenStudy (jackellyn):

Right, now solve for x

OpenStudy (anonymous):

x=4

OpenStudy (jackellyn):

Show me your work please?

OpenStudy (anonymous):

-x+-10=-6-10 subtract -10 from both sides x=4

OpenStudy (jackellyn):

You add 10 to both sides -x-10+10=-6-10 -x=4 x=-4

OpenStudy (anonymous):

oh ok thank u

OpenStudy (jackellyn):

Now you can plot (-4,-5) on a graph